What is the final value of the system described by the transfer function: if the input is 10u(t)?
Answer #1

# [+] = loult), Apply LiT X15) = 10 475) = H (S). X6) = st2 s+45+2 10 1 46) = 10(5+2) 545745+23 final value y (00) = It sys)